OpenStudy (anonymous):

Hi guys, I need help with a sequence problem : Find the next term in this sequence : 24,54,90,120,168,210,240,288,360,440 I've been trying this for quite some time now and I'm not getting the logic. Help is solicited. Thanks in advance!

JoelTheBoss (joel_the_boss):

Hello Welcome to OpenStudy! Can you tell which part you are stuck at exactly so I can help you. Thank you. :)

JoelTheBoss (joel_the_boss):

30 36 30 48 42 30 48 72 80 Those would be the changes in between each of the numbers. Does that clear it up a little? Are there any options for this question? @ripazha

OpenStudy (anonymous):

thanks for the reply! i am trying to find the pattern here. I calculated the term differences. That turned out to be non-constructive. I don't know what to do after that

OpenStudy (anonymous):

\(\rlap{\color{red}{\huge\bigstar}\huge \color{green}{ \text{Welcome to Open Study! }}\color{red}\bigstar}{\; \color{aqua}{\huge\bigstar}\huge \color{aqua}{\text{Welcome to Open Study! }}\color{yellow}\bigstar}\) @ripazha

OpenStudy (anonymous):

I don't see the pattern from the differences. can you elaborate? thanks

OpenStudy (michele_laino):

all of your numbers are divisible by 6, except the last, namely 440

hartnn (hartnn):

if there's no pattern in differences, then the sequence is not a polynomial sequence..

OpenStudy (anonymous):

@hartnn , so? What should I do after that?

JoelTheBoss (joel_the_boss):

Are you learning about Cubed numbers? Or squared numbers? Or Triangular number? Or Geometric sequences?

OpenStudy (anonymous):

this was in one of the iq tests at school

OpenStudy (michele_laino):

is it 990?

JoelTheBoss (joel_the_boss):

To find a missing number, we need the rule. Is there something that I missed? Or is there no rule yet?

hartnn (hartnn):

we need to find that rule/pattern joel

OpenStudy (anonymous):

@MARC_ How?

hartnn (hartnn):

any choices are given ?

JoelTheBoss (joel_the_boss):

FOUND IT! I think.. It ha only worked with 2 of them so far give me a sec...

OpenStudy (anonymous):

@Michele_Laino , How?

OpenStudy (ganpat):

24 = 6 * 4 54 = 6 * 9 90 = 6 * 15 120 = 6 * 20 168 = 6 * 28 210 = 6 * 35 240 = 6 * 40 288 = 6 * 48 360 = 6 * 60 440 = 6 * 73.33 i guess the question is wrong...

OpenStudy (anonymous):

It should be around 520

hartnn (hartnn):

ganpat, but there is no pattern in 4,9,15,20....is it ?

JoelTheBoss (joel_the_boss):

hmm ever mind I thought of it wrong...

OpenStudy (anonymous):

Guys, I appreciate your help but it would really help if you mention the logic

OpenStudy (anonymous):

\(2^3 . 3\) \(2 .3^2\) \(2 . 3^2.5\) \(2^3. 3.5\) \(2^3. 3.7\) \(2 . 3.5.7\) \(2^2. 3.5.7\) \(2^5. 3^2 \) \(2^3. 3^2.5\) \(3^2. 5.11\)

OpenStudy (ganpat):

@hartnn : right.. so i feel the question is wrong or may some values are wrongly written.

OpenStudy (anonymous):

so since 3 every where lets see what we got

JoelTheBoss (joel_the_boss):

Ok first off before we jump to an answer., It would be nice to know what you learning about. Are you in geometry or algebra, or what?

OpenStudy (anonymous):

@Ganpat question is right. my teacher was confident and he has a solution which he won't reveal

OpenStudy (anonymous):

wait i have typo

OpenStudy (anonymous):

\(2^3 . 3\) \(2 .3^2\) \(2 . 3^2.5\) \(2^3. 3.5\) \(2^3. 3.7\) \(2 . 3.5.7\) \(2^2. 3.5.7\) \(2^5. 3^2 \) \(2^3. 3^2.5\) \(2^3. 5.11\)

hartnn (hartnn):

one more 2nd line \(2.3^3\)

OpenStudy (anonymous):

so 2 every where not 3 :P

OpenStudy (anonymous):

\(2^3 . 3\) \(2 .3^3\) \(2 . 3^2.5\) \(2^3. 3.5\) \(2^3. 3.7\) \(2 . 3.5.7\) \(2^2. 3.5.7\) \(2^5. 3^2 \) \(2^3. 3^2.5\) \(2^3. 5.11\) now fine ?

OpenStudy (michele_laino):

since your sequence is: 24=6*4 54=6*9 90=6*15 .... 360=6*60 now 440 is not divisible by 6, nevertheless if we note that: 440=110*4 now I start a sequence as above, with 110 in plce of 6 namely the next term is: 990=110 *9

OpenStudy (ganpat):

@ripazha : ha haa.. No offenses.. just a second thought :p Rule of Mathematics: If you can`t prove it right, then prove it wrong :D

OpenStudy (michele_laino):

oops ...in place of...

OpenStudy (anonymous):

I give up... whoever solves this is a genius

OpenStudy (anonymous):

@Michele_Laino , according to your argument, where did 440 come from? Your argument doesn't explain the presence of 440

OpenStudy (mathmath333):

can u post a screenshot of question

OpenStudy (anonymous):

@mathmath333 ,why?

OpenStudy (mathmath333):

i will put that in a supercomputer

OpenStudy (mathmath333):

kidding , just want to know if anything is incorrect or misspelled

OpenStudy (anonymous):

@mathmath333 I'm 100 % sure that the question is right. I've checked a dozen times

OpenStudy (mathmath333):

cuz yesterday o guy posted \(9\) as \(g\)

OpenStudy (michele_laino):

I thought that was an arbitrary change in your sequence @ripazha

OpenStudy (anonymous):

@Michele_Laino I don't think so. @mathmath333 again. I'm 100% sure about the question

OpenStudy (mathmath333):

yes , but thats the only chance if any i think

OpenStudy (anonymous):

@mathmath333 ?

OpenStudy (anonymous):

i got weird pattern

hartnn (hartnn):

for now even weird is good enough to consider :P

OpenStudy (mrnood):

Ok - I have plotted these numbers. their differences and their sin It ia fairly evident that there is no arithmetic, geometric or trig relationship between them. You said that this came from an IQ test. Now it is my experience that IQ tests are not always mathematically related, and sometimes require a you to spot a relationship other than purely mathematical. Right now I don't see a pattern - but I think we may be looking in the wrong place if we pursue a rigourous mathematical approach. Nice question...

hartnn (hartnn):

i guess we need to think out of the box!

OpenStudy (anonymous):

n x(n) div by 2 x(n)/2 - x(n-1)/2 x(n)/2 mod x(n-1)/2 1 24 12 Null Null 2 54 27 15 3 or 15 3 90 45 18 18 4 120 60 15 15 5 168 84 24 24 6 210 105 21 21 7 240 120 15 15 8 288 144 24 24 9 360 180 36 36 10 440 220 40 40

OpenStudy (anonymous):

so nw we can find something :D

OpenStudy (anonymous):

lets try it :- x(1)=2*12 lets find x(2) x(2)/2-12=x(2)/2 mod 12 so x(2)/2=27 x(2)=54

OpenStudy (anonymous):

which is something silly and i just note it wrong >.<

OpenStudy (anonymous):

lol

OpenStudy (michele_laino):

is it a number close to 450?

OpenStudy (anonymous):

@Michele_Laino , as i've mentioned before, I don't know the answer. If you give the logic, it'd be great

OpenStudy (mrnood):

@ripazha Perhaps it is NOT logic. Look at the numbers, maybe there is a link to non-sequential items. scores in games? number of days between moonshots, number of months each Pope lived ..... blah blah - all I'm saying is that for IQ tests you cannot rely on maths - you may need to be more creative.

OpenStudy (anonymous):

@MrNood , I don't think so. My teacher said there is A solution.

OpenStudy (michele_laino):

ok! I write your sequence: 6*4-6*9-6*15-6*20-6*28-6*35-6*40-6*48-6*60-440 now (4+15)/2-1=9, so the mid term between 24 and 60, is 6*9=54, at the same way, if I suppose 444=6*74 instead of 440, I have: (48+74)/2-1=60, so the mid term between 288 and 440 is 6*60=360. Again for the next term, namely y=6*x we have: (x+60)/2-1=74, from which x=90, so the next term to 444 is 6*90=450.

OpenStudy (michele_laino):

and that reasoning would explain the presence of 440

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

it is nto there in oeis and wolfram also gave up so im not so sure.. id like to know the source of this problem / level of difficulty before attempting

ganeshie8 (ganeshie8):

it is school level iq test so it should be simpler than mods and congruences stuff hmm

OpenStudy (anonymous):

i know it should be clear observation , but see not linear relation neither cubic or quadratic , was thinking it should be something without calculating like digits replacement

ganeshie8 (ganeshie8):

yeah im trying interleaving sequences and other non-math related things it cannot be a polynomial because the differences are not converging and also it cannot be an exponential because the differences keep increasign and decreasign

OpenStudy (anonymous):

i agree

ganeshie8 (ganeshie8):

no success :O il wait for the teacher to show how to solve i guess :)

OpenStudy (anonymous):

Hey guys, I called my teacher to ask about this, and he said ; Do whatever you want. I just want a closed form solution. You are free to use any operations.